# Less is More: Data-Efficient Adaptation for Controllable Text-to-Video Generation

### v2

@@ 1 Introduction
Recent advances in generative AI using diffusion models have enabled unprecedented levels of quality in video generation The primary foundational models for video generation are text-to-video (T2V) such as [wan2025wan, polyak2024movie], where users describe their desired creations in natural language. Due to limitations of controllability via text, significant effort has been put into creating methods that accept other input modalities, such as images, keyframes, depth maps, bounding boxes, pose skeletons, driver videos, camera trajectories, and so on [wan2025wan, jiang2025vace, cheng2025wan, bai2025recammaster]. However, achieving consistent, reliable, and intuitive fine-grained control over all aspects of the output video still remains a challenge. In this work, we tackle the problem of adding a control mechanism over specific low-dimensional physical or optical properties, such as camera intrinsics, via a simple synthetic data generation and model fine-tuning framework.
@@
A common pattern that has emerged in methods for producing specialized generative video models is starting with a large “foundation” model, trained on huge amounts of video data, and fine-tuning it on a carefully crafted task-specific smaller dataset [polyak2024movie]. Such datasets help the model focus on a particular character identity, artistic style, or specialized effect. The success of such approaches hints at the fact that the initial pre-training equips the model with many useful priors implicit in its latent representation, which can be explicitly “coaxed” out during post-training. Our approach is the first approach that aims to enable conditioning on camera effects (like shutter speed, focal length, etc) in pre-trained video generative models for consistent generation. While the quality of the post-training data is indeed critical, we argue and demonstrate that, surprisingly, having data that are perfectly photorealistic and representative of the output domain can not only be unnecessary but even detrimental for certain specializations.
@@
This paper proposes an approach that allows the use of synthetic data to learn several conditioning effects. We contribute a novel joint training approach that factorizes adaptation: a standard low-rank adapter (LoRA) [hu2022lora] encodes a minimal domain shift, while a disentangled cross-attention adapter learns the conditioning physical effect. Additionally, our paper’s contributions include a formal analysis of why this data-efficient approach succeeds. We demonstrate that, counterintuitively, the success of small-data fine-tuning hinges on the simplicity of the synthetic data. We show that fine-tuning on photorealistic synthetic data, while seemingly a higher-fidelity choice, induces catastrophic forgetting by corrupting the backbone’s pre-trained priors, leading to a “content collapse.” To quantify this, we introduce a new evaluation framework that measures this generative drift and its impact on semantic fidelity. Using this framework, we show that a model trained on simple data retains its generative diversity and high semantic fidelity, while the model trained on complex data suffers a quantifiable and catastrophic collapse. Our work provides both a data-efficient method for controllable video generation and a formal methodology for diagnosing and preventing backbone corruption during adaptation.
@@ 2 Related Work
Text-to-Video Generation. Diffusion models [ho2020denoising, sohl2015deep] have become the leading approach for video generation, largely by adapting pre-trained text-to-image models with temporal modules. Early methods like AnimateDiff [guo2023animatediff] showed that adding temporal attention to a frozen T2I U-Net [ronneberger2015u] enables high-quality motion. The field has since moved toward Diffusion Transformers (DiTs) [peebles2023scalable], which underpin state-of-the-art open-source systems such as Hunyuan Video [kong2024hunyuanvideo] and Wan [wan2025wan]. Closed-source models including Sora [OpenAI_Sora_2024], Gen-3 [Runway_Gen3_2024], and Kling [Kuaishou_Kling_2024] further push the frontier with long-form, high-fidelity, physically consistent video. Although our experiments use Wan T2V model, our fine-tuning paradigm and evaluation framework are model-agnostic and apply to any video diffusion backbone.
@@
Controllable Generative Models. Recent work introduces explicit control in generative models through spatial conditions such as depth, masks, sketches, boxes, and segmentation maps [bhat2024loosecontrol, wang2023imagen, voynov2023sketch, li2023gligen, couairon2022diffedit]. ControlNet [zhang2023adding] and T2I-Adapters [mou2024t2i] support such inputs efficiently, with extensions to video focusing on camera pose and trajectory [he2024cameractrl, yu2024viewcrafter, wang2024motionctrl]. Methods like Bokeh Diffusion [fortes2025bokeh] and Generative Photography [Yuan_2024_GenPhoto] model physical imaging effects, but their image-based designs fail to generalize to video due to temporal inconsistency. Token-based conditioning via CSaT [fang2024camera] offers implicit control but remains data-heavy. Large-scale, richly annotated video datasets for physical camera parameters are challenging to obtain. Image LoRAs [hu2022lora] provide lightweight adaptation for effects including motion blur, bokeh, restoration, and super-resolution [zhang2025lion, fortes2025bokeh, park2024contribution, sun2025pixel]. Our approach builds on these ideas by injecting scalar physical parameters through a decoupled cross-attention adapter for controllable video synthesis, enabled by targeted synthetic data.
@@
Learning from Synthetic and Sparse Data. The difficulty of large-scale video collection has led to work showing that strong representations can emerge from simple synthetic data [tian2024learning, mayer2018makes]. Procedurally generated videos can rival natural-video pretraining [yu2024learning], and synthetic datasets often serve as effective substitutes for real videos [guo2022learning, kim2022transferable]. Synthetic 3D environments also provide precise labels for tasks like optical flow, tracking, and video depth [dosovitskiy2015flownet, zheng2023pointodyssey, yang2024depth], and recent methods use such data for low-shot motion learning [ren2024customize, wu2024lamp]. These results suggest that motion priors need not rely on complex natural data, motivating our use of sparse, task-specific synthetic supervision to condition T2V models on camera parameters such as shutter speed, aperture, and temperature.
@@
Evaluation and Metrics for Generative Video Existing one-shot fine-tuning methods [wu2023tune] and even zero-shot approaches [geyer2023tokenflow, qi2023fatezero, yang2023rerender] still struggle to generate high-quality, varied videos. Even just quantitatively evaluating and comparing against such such models remains a complex, multi-dimensional problem. Common metrics include the Fréchet Video Distance (FVD) [unterthiner2018towards, unterthiner2019fvd], which measures distributional similarity in a learned feature space for action recognition, and CLIP Score [radford2021learning], which quantifies text–video semantic alignment. More recent works have proposed adaptations for video, such as X-CLIP Score [ni2022expanding], or more robust alternatives like VQA Score [lin2024evaluating], which is less prone to error on complex compositional prompts. shuttleworth2025loravsfinetuningillusion further investigate how LoRA and full fine-tuning alter pre-trained diffusion models by analyzing changes in their weight matrices through spectral decomposition. While benchmarks such as VBench [huang2024vbench]—which we adopt in our evaluation—offer a suite of metrics for assessing temporal consistency, object fidelity, and aesthetic quality, we identify a critical methodological gap: standard metrics fail to quantify the intrinsic complexity of the fine-tuning data. We address this with a novel metric as part of our contribution.
@@ 3 Method
Disentangled Conditioning Module Following recent work in controllable generation [ye2023ip], we employ a dedicated cross-attention module in order to inject the conditioning signal. We normalize the condition (which in our experiments is a scalar value) to obtain $c\in[-1,1]$. This condition is projected to a higher-dimension embedding vector via a small multi-layer perceptron (MLP): $e_{\text{cond}}=\operatorname{MLP}_{\text{cond}}(c)$. The embedded conditional signal $e_{cond}$ is then injected into the model using a dedicated cross-attention mechanism that operates in parallel with the main text cross-attention. We insert this conditional cross-attention adapter into the deepest third of the original model’s transformer blocks. For a given query $q$ from the video latent, the output of the cross-attention block is a linear combination of the text-conditioning $y_{\text{text}}$ and the attribute-conditioning $y_{\text{cond}}$. Backbone LoRA Rather than attempting to construct a dataset for fine-tuning that uniformly and consistently exhibits variation along the desired control axis and matches the natural video distribution of the original T2V model, we make the deliberate choice to construct datasets that satisfy only the former requirement. This decision frees us from having to do careful curation of natural videos or difficult photorealistic rendering but rather allows us to focus on creating simplistic synthetic examples that illustrate the conditioning effect. However, fine-tuning on such out-of-domain datasets inevitably induces content drift. To address this, we employ a joint training strategy where a backbone LoRA is optimized alongside the conditioning module. This approach enforces a “separation of concerns”: the backbone LoRA absorbs the synthetic domain shift, allowing the conditional adapter to focus exclusively on disentangling the physical control signal. Critically, at inference time, we revert this drift by selectively discarding the backbone LoRA weights from all blocks that are not equipped with the conditional adapter. This restores the model’s original generative priors in the majority of the network while preserving the learned control mechanism.
@@ 4 Dataset Construction
  • Aperture: We generate depth-of-field variations by applying BokehMe  [Peng2022BokehMe] to a single image featuring a complex foreground subject and detailed background.
@@ 5.2 Stage 2: Slow Validation Protocol (SVP)
  • Semantic Fidelity (Backbone Health): We compute video-text alignment using the X-CLIP Score [ni2022expanding] and the VQA score [lin2024evaluating]. Cross-referencing these scores with the $\mathcal{V}_{\text{drift}}$ from FEP allows us to diagnose the adaptation: a high-velocity drift is benign if final scores are high (successful adaptation), but malignant if they collapse (catastrophic forgetting).
  • Video Quality & Temporal Coherence: We use six metrics from VBench [zhang2024evaluationagent] to evaluate the generated video quality independent of text guidance: subject/background consistency, motion smoothness, dynamic degree, aesthetic quality, and imaging quality.
@@ 6 Experiments
To validate our “Less is More” hypothesis and analyze how different data strategies impact backbone fidelity, we conduct two primary groups of experiments. We use the WAN 2.1 [wan2025wan] as our T2V backbone for all fine-tuning and evaluation.
@@ 6.1 Experimental Design
  • Group 1: Data Complexity (Syn vs. Real). To isolate the effect of data content, we conduct a controlled one-shot comparison. For each camera parameter, we train two models on minimal data: (1) a model trained on a single synthetic scene and (2) a model trained on a single photorealistic scene (per Sec. 4). From each scene, we sample 7 scalar conditions to create the training data. This ablation is evaluated using our FEP on the 800-prompt VBench benchmark [zhang2024evaluationagent].
@@ 6.3 Visual Condition Comparison
Apart from the comparisons via training data and inference methods, we also conducted some qualitative comparisons to prior works such as WAN 2.1 [wan2025wan], Bokeh Diffusion [fortes2025bokeh], and [Yuan_2024_GenPhoto]. Fig. 5 compares the condition ability of different workflow, text-based prompting in the T2V backbone (WAN 2.1) exhibits significant semantic entanglement. This is particularly evident in the Shutter Speed test (Row 1), where the baseline either struggles to associate shutter speed changes with motion blur, or generates a lower-fidelity scene by creating a “metalic” door instead of the requested glass one. A similar entanglement occurs in the Temperature test (Row 3), where the baseline misinterprets color temperature as weather, incorrectly altering the scene’s environment (e.g., adding snow) rather than its color tone. Our method overcomes these shortcomings by correctly applying the physical effect of shutter speed (motion blur) without this semantic or fidelity confusion. Furthermore, for Aperture (Row 2), our model generates high-fidelity bokeh effects comparable to specialized, data-intensive baselines such as Bokeh Diffusion [fortes2025bokeh] and Generative Photography [Yuan_2024_GenPhoto]. These results validate our central hypothesis: our data-efficient synthetic training achieves robust physical disentanglement via a low-rank adapter, eluding the limitations of entangled text prompts and rivaling the quality of data-heavy methods. The supplementary material (see Appendix F) provides a number of additional examples (as videos) of our method that demonstrate the consistency of on a range of control values.
Figure 5: Qualitative comparison of physical controls. We compare our method against text-based prompting in a T2V backbone (WAN 2.1) and specialized, data-intensive image baselines [Yuan_2024_GenPhoto] and Bokeh Diffusion [fortes2025bokeh].
@@ 7.1 Hypothesis 1: Low-Fidelity Data Prevents Catastrophic Forgetting
Method: Backbone Intruder Dimension Check. We apply the “intruder dimension” analysis from  shuttleworth2024lora. We perform an SVD on the original backbone weight $W_{pre}$ and its final adapted version $W_{lora}=W_{pre}+\Delta W_{lora}$.
@@ Appendix C Training Details
We use four condition-adapter tokens, an adapter embedding dimension of 256, and LoRA rank 32 for both the condition adapters and the backbone LoRA, with the adapter gate fixed at 0.5 throughout training. A learning rate of $2{\times}10^{-5}$ with a 100-step warmup is applied. Optimization is performed using the AdamW optimizer[loshchilov2017decoupled] with betas $(0.9,0.99)$, weight decay 0.01, and $\varepsilon=10^{-8}$. All LoRA and adapter parameters are trained in bfloat16 with activation checkpointing enabled for memory efficiency.

### v3

@@ 1 Introduction
Recent advances in generative AI using diffusion models have enabled unprecedented levels of quality in video generation The primary foundational models for video generation are text-to-video (T2V) such as [43, 28], where users describe their desired creations in natural language. Due to limitations of controllability via text, significant effort has been put into creating methods that accept other input modalities, such as images, keyframes, depth maps, bounding boxes, pose skeletons, driver videos, camera trajectories, and so on [43, 15, 3, 1]. However, achieving consistent, reliable, and intuitive fine-grained control over all aspects of the output video still remains a challenge. In this work, we tackle the problem of adding a control mechanism over specific low-dimensional physical or optical properties, such as camera intrinsics, via a simple synthetic data generation and model fine-tuning framework.
@@
A common pattern that has emerged in methods for producing specialized generative video models is starting with a large “foundation” model, trained on huge amounts of video data, and fine-tuning it on a carefully crafted task-specific smaller dataset [28]. Such datasets help the model focus on a particular character identity, artistic style, or specialized effect. The success of such approaches hints at the fact that the initial pre-training equips the model with many useful priors implicit in its latent representation, which can be explicitly “coaxed” out during post-training. Our approach is the first approach that aims to enable conditioning on camera effects (like shutter speed, focal length, etc) in pre-trained video generative models for consistent generation. While the quality of the post-training data is indeed critical, we argue and demonstrate that, surprisingly, having data that are perfectly photorealistic and representative of the output domain can not only be unnecessary but even detrimental for certain specializations.
@@
This paper proposes an approach that allows the use of synthetic data to learn several conditioning effects. We contribute a novel joint training approach that factorizes adaptation: a standard low-rank adapter (LoRA) [13] encodes a minimal domain shift, while a disentangled cross-attention adapter learns the conditioning physical effect. Additionally, our paper’s contributions include a formal analysis of why this data-efficient approach succeeds. We demonstrate that, counterintuitively, the success of small-data fine-tuning hinges on the simplicity of the synthetic data. We show that fine-tuning on photorealistic synthetic data, while seemingly a higher-fidelity choice, induces catastrophic forgetting by corrupting the backbone’s pre-trained priors, leading to a “content collapse.” To quantify this, we introduce a new evaluation framework that measures this generative drift and its impact on semantic fidelity. Using this framework, we show that a model trained on simple data retains its generative diversity and high semantic fidelity, while the model trained on complex data suffers a quantifiable and catastrophic collapse. Our work provides both a data-efficient method for controllable video generation and a formal methodology for diagnosing and preventing backbone corruption during adaptation.
@@ 2 Related Work
Text-to-Video Generation. Diffusion models [12, 36] have become the leading approach for video generation, largely by adapting pre-trained text-to-image models with temporal modules. Early methods like AnimateDiff [10] showed that adding temporal attention to a frozen T2I U-Net [32] enables high-quality motion. The field has since moved toward Diffusion Transformers (DiTs) [26], which underpin state-of-the-art open-source systems such as Hunyuan Video [17] and Wan [43]. Closed-source models including Sora [24], Gen-3 [33], and Kling [38] further push the frontier with long-form, high-fidelity, physically consistent video. Although our experiments use Wan T2V model, our fine-tuning paradigm and evaluation framework are model-agnostic and apply to any video diffusion backbone.
@@
Controllable Generative Models. Recent work introduces explicit control in generative models through spatial conditions such as depth, masks, sketches, boxes, and segmentation maps [2, 44, 42, 18, 4]. ControlNet [55] and T2I-Adapters [22] support such inputs efficiently, with extensions to video focusing on camera pose and trajectory [11, 51, 45]. Methods like Bokeh Diffusion [7] and Generative Photography [53] model physical imaging effects, but their image-based designs fail to generalize to video due to temporal inconsistency. Token-based conditioning via CSaT [6] offers implicit control but remains data-heavy. Large-scale, richly annotated video datasets for physical camera parameters are challenging to obtain. Image LoRAs [13] provide lightweight adaptation for effects including motion blur, bokeh, restoration, and super-resolution [56, 7, 25, 37]. Our approach builds on these ideas by injecting scalar physical parameters through a decoupled cross-attention adapter for controllable video synthesis, enabled by targeted synthetic data.
@@
Learning from Synthetic and Sparse Data. The difficulty of large-scale video collection has led to work showing that strong representations can emerge from simple synthetic data [39, 21]. Procedurally generated videos can rival natural-video pretraining [52], and synthetic datasets often serve as effective substitutes for real videos [9, 16]. Synthetic 3D environments also provide precise labels for tasks like optical flow, tracking, and video depth [5, 57, 48], and recent methods use such data for low-shot motion learning [31, 47]. These results suggest that motion priors need not rely on complex natural data, motivating our use of sparse, task-specific synthetic supervision to condition T2V models on camera parameters such as shutter speed, aperture, and temperature.
@@
Evaluation and Metrics for Generative Video Existing one-shot fine-tuning methods [46] and even zero-shot approaches [8, 29, 49] still struggle to generate high-quality, varied videos. Even just quantitatively evaluating and comparing against such such models remains a complex, multi-dimensional problem. Common metrics include the Fréchet Video Distance (FVD) [40, 41], which measures distributional similarity in a learned feature space for action recognition, and CLIP Score [30], which quantifies text–video semantic alignment. More recent works have proposed adaptations for video, such as X-CLIP Score [23], or more robust alternatives like VQA Score [19], which is less prone to error on complex compositional prompts. Shuttleworth et al. [35] further investigate how LoRA and full fine-tuning alter pre-trained diffusion models by analyzing changes in their weight matrices through spectral decomposition. While benchmarks such as VBench [14]—which we adopt in our evaluation—offer a suite of metrics for assessing temporal consistency, object fidelity, and aesthetic quality, we identify a critical methodological gap: standard metrics fail to quantify the intrinsic complexity of the fine-tuning data. We address this with a novel metric as part of our contribution.
@@ 3 Method
Disentangled Conditioning Module Following recent work in controllable generation [50], we employ a dedicated cross-attention module in order to inject the conditioning signal. We normalize the condition (which in our experiments is a scalar value) to obtain $c\in[-1,1]$. This condition is projected to a higher-dimension embedding vector via a small multi-layer perceptron (MLP): $e_{\text{cond}}=\operatorname{MLP}_{\text{cond}}(c)$. The embedded conditional signal $e_{cond}$ is then injected into the model using a dedicated cross-attention mechanism that operates in parallel with the main text cross-attention. We insert this conditional cross-attention adapter into the deepest third of the original model’s transformer blocks. For a given query $q$ from the video latent, the output of the cross-attention block is a linear combination of the text-conditioning $y_{\text{text}}$ and the attribute-conditioning $y_{\text{cond}}$. Backbone LoRA Rather than attempting to construct a dataset for fine-tuning that uniformly and consistently exhibits variation along the desired control axis and matches the natural video distribution of the original T2V model, we make the deliberate choice to construct datasets that satisfy only the former requirement. This decision frees us from having to do careful curation of natural videos or difficult photorealistic rendering but rather allows us to focus on creating simplistic synthetic examples that illustrate the conditioning effect. However, fine-tuning on such out-of-domain datasets inevitably induces content drift. To address this, we employ a joint training strategy where a backbone LoRA is optimized alongside the conditioning module. This approach enforces a “separation of concerns”: the backbone LoRA absorbs the synthetic domain shift, allowing the conditional adapter to focus exclusively on disentangling the physical control signal. Critically, at inference time, we revert this drift by selectively discarding the backbone LoRA weights from all blocks that are not equipped with the conditional adapter. This restores the model’s original generative priors in the majority of the network while preserving the learned control mechanism.
@@ 4 Dataset Construction
  • Aperture: We generate depth-of-field variations by applying BokehMe  [27] to a single image featuring a complex foreground subject and detailed background.
@@ 5.2 Stage 2: Slow Validation Protocol (SVP)
  • Semantic Fidelity (Backbone Health): We compute video-text alignment using the X-CLIP Score [23] and the VQA score [19]. Cross-referencing these scores with the $\mathcal{V}_{\text{drift}}$ from FEP allows us to diagnose the adaptation: a high-velocity drift is benign if final scores are high (successful adaptation), but malignant if they collapse (catastrophic forgetting).
  • Video Quality & Temporal Coherence: We use six metrics from VBench [54] to evaluate the generated video quality independent of text guidance: subject/background consistency, motion smoothness, dynamic degree, aesthetic quality, and imaging quality.
@@ 6 Experiments
To validate our “Less is More” hypothesis and analyze how different data strategies impact backbone fidelity, we conduct two primary groups of experiments. We use the WAN 2.1 [43] as our T2V backbone for all fine-tuning and evaluation.
@@ 6.1 Experimental Design
  • Group 1: Data Complexity (Syn vs. Real). To isolate the effect of data content, we conduct a controlled one-shot comparison. For each camera parameter, we train two models on minimal data: (1) a model trained on a single synthetic scene and (2) a model trained on a single photorealistic scene (per Sec. 4). From each scene, we sample 7 scalar conditions to create the training data. This ablation is evaluated using our FEP on the 800-prompt VBench benchmark [54].
@@ 6.3 Visual Condition Comparison
Apart from the comparisons via training data and inference methods, we also conducted some qualitative comparisons to prior works such as WAN 2.1 [43], Bokeh Diffusion [7], and [53]. Fig. 5 compares the condition ability of different workflow, text-based prompting in the T2V backbone (WAN 2.1) exhibits significant semantic entanglement. This is particularly evident in the Shutter Speed test (Row 1), where the baseline either struggles to associate shutter speed changes with motion blur, or generates a lower-fidelity scene by creating a “metalic” door instead of the requested glass one. A similar entanglement occurs in the Temperature test (Row 3), where the baseline misinterprets color temperature as weather, incorrectly altering the scene’s environment (e.g., adding snow) rather than its color tone. Our method overcomes these shortcomings by correctly applying the physical effect of shutter speed (motion blur) without this semantic or fidelity confusion. Furthermore, for Aperture (Row 2), our model generates high-fidelity bokeh effects comparable to specialized, data-intensive baselines such as Bokeh Diffusion [7] and Generative Photography [53]. These results validate our central hypothesis: our data-efficient synthetic training achieves robust physical disentanglement via a low-rank adapter, eluding the limitations of entangled text prompts and rivaling the quality of data-heavy methods. The supplementary material (see Appendix F) provides a number of additional examples (as videos) of our method that demonstrate the consistency of on a range of control values.
Figure 5: Qualitative comparison of physical controls. We compare our method against text-based prompting in a T2V backbone (WAN 2.1) and specialized, data-intensive image baselines [53] and Bokeh Diffusion [7].
@@ 7.1 Hypothesis 1: Low-Fidelity Data Prevents Catastrophic Forgetting
Method: Backbone Intruder Dimension Check. We apply the “intruder dimension” analysis from  Shuttleworth et al. [34]. We perform an SVD on the original backbone weight $W_{pre}$ and its final adapted version $W_{lora}=W_{pre}+\Delta W_{lora}$.
@@ Appendix C Training Details
We use four condition-adapter tokens, an adapter embedding dimension of 256, and LoRA rank 32 for both the condition adapters and the backbone LoRA, with the adapter gate fixed at 0.5 throughout training. A learning rate of $2{\times}10^{-5}$ with a 100-step warmup is applied. Optimization is performed using the AdamW optimizer[20] with betas $(0.9,0.99)$, weight decay 0.01, and $\varepsilon=10^{-8}$. All LoRA and adapter parameters are trained in bfloat16 with activation checkpointing enabled for memory efficiency.